# Optimization of Superconducting Niobium Nitride Thin Films via High-Power Impulse Magnetron Sputtering

### v1

@@ Abstract
We report a systematic comparison of niobium nitride thin films deposited on oxidized silicon substrates by reactive DC magnetron sputtering and reactive high-power impulse magnetron sputtering (HiPIMS). After determining the nitrogen gas concentration that produces the highest superconducting critical temperature for each process, we characterize the dependence of the critical temperature on film thickness. The optimal nitrogen concentration is higher for HiPIMS than for DC sputtering, and HiPIMS produces higher critical temperatures for all thicknesses studied. We attribute this to the HiPIMS process enabling the films to get closer to optimal stoichiometry before beginning to form a hexagonal crystal phase that reduces the critical temperature, along with the extra kinetic energy in the HiPIMS process enabling greater adatom mobility and improving crystallinity. We also study the effects of an aluminum nitride buffer layer and substrate heating on the critical temperature.
@@ I Introduction
Niobium nitride (NbN) is a refractory metal with a superconducting critical temperature ($T_{\mathrm{C}}$) as high as $\approx 16$ K.Shy, Toth, and Somasundaram (1973) NbN thin films are utilized in a wide range of superconducting device applications, including photodetectors such as hot electron bolometersGousev et al. (1994) and superconducting nanowire single-photon detectors;Holzman and Ivry (2019) and Josephson junction devices such as SIS heterodyne mixers,McGrath et al. (1991) RSFQ digital logic,Shao et al. (2024) and superconducting qubits.Kim et al. (2021) Due to their high kinetic inductivity, NbN thin films are of interest in applications such as slow-wave transmission lines,Santavicca et al. (2021) tunable couplers,Colangelo et al. (2021) and travelling wave parametric amplifiers.Adamyan et al. (2016) For many of these applications, very thin films ($\lesssim 10$ nm) are required, and hence a number of groups have explored different techniques for producing high-quality ultra-thin NbN films.Ilin et al. (2008); Espiau de Lamaëstre, Odier, and Villégier (2007); Iovan et al. (2023); Chockalingam et al. (2008); Wright, Xing, and Jena (2023); Schneider et al. (2009); Shiino et al. (2010); Rhazi et al. (2021)
Several crystal phases of NbN can occur under normal conditions. The cubic NaCl-like $\delta$-phase has the highest $T_{\mathrm{C}}$ and occurs when a near-stoichiometric reaction between nitrogen and niobium takes place.Shy, Toth, and Somasundaram (1973) Increasing the nitrogen to a $1{:}1$ stoichiometry can result in the formation of the hexagonal $\epsilon$-phase with a lower $T_{\mathrm{C}}$. On the other hand, reducing the nitrogen concentration can lead to the formation of tetragonal Nb4N3 or hexagonal Nb2N, which also have a lower $T_{\mathrm{C}}$.Benkahoul et al. (2004); Wright, Xing, and Jena (2023) Optimizing the nitrogen concentration is thus important for maximizing the $T_{\mathrm{C}}$ of NbN thin films.
@@
Heating the substrate can improve the crystallinity of the film, and single-domain epitaxial films can be produced using an appropriate substrate at sufficiently high temperature.Kim et al. (2021); Wright, Xing, and Jena (2023); Iovan et al. (2023); Chockalingam et al. (2008) These high temperatures, however, are not compatible with some materials and fabrication processes, including lift-off processes and standard CMOS processes. Single-domain epitaxial films also have a lower kinetic inductivity than polycrystaline films and hence may not be desirable for applications that utilize the high kinetic inductivity of NbN.
@@ II Deposition Parameters
NbN films were deposited using a Nb target in a reactive process with nitrogen gas. The Nb target was specified as 99.95% pure (excluding tantalum). Ultra high purity argon gas (99.9999% purity) and ultra high purity nitrogen gas (99.999% purity) were introduced into the chamber via separate mass flow controllers. The chamber was pumped to $\lesssim 5\times 10^{-7}$ torr prior to deposition. During all depositions, a gate valve is used to throttle the turbomolecular pump and the mass flow controllers set the flow rates to achieve a pressure of $3.0\times 10^{-3}$ torr. The nitrogen gas concentration is specified as a percentage of the argon gas flow rate (not as a percentage of the total gas flow). For example, an argon flow rate of $12.5$ sccm and a nitrogen flow rate of $2.5$ sccm would be specified as a $20$% nitrogen gas concentration.
Depositions were performed using both DC sputtering and high-power impulse magnetron sputtering (HiPIMS). The HiPIMS depositions utilized an Impulse $2$ kW pulsed power module from Starfire Industries. This module allows for a user-configurable voltage pulse width, height, and repetition rate. All HiPIMS depositions described here used a $1$ kHz repetition rate and a $100$ $\mu$s pulse width. The HiPIMS was performed in a power-limited mode set to a time-average power of $250$ W; this determined the resulting voltage pulse height, which had a value of approximately $480$ V. The main pulse was followed by a positive kick pulse with an amplitude of approximately $+50$ V and a duration of $50$ $\mu$s. This kick pulse has been shown to increase the deposition rate by driving ionized target material to the sample, as the high peak power of the HiPIMS process can result in a higher ionization fraction of the target material.Gaines (2019); Kalal et al. (2022) An example of the resulting voltage and current pulses is shown in figure 1. The peak power, which occurs at the maximum of the current pulse, is approximately $3$ kW.
@@ III Comparison of DC and HiPIMS Films
NbN films were deposited onto unheated substrates at varying nitrogen gas concentrations using both DC sputtering and HiPIMS. The deposition rate varies with the nitrogen gas concentration, so the deposition rate had to be determined for each nitrogen concentration. This was done by performing a $900$ s deposition, scribing and snapping the resulting sample, and imaging the sample’s cross-section in a scanning electron microscope (SEM) to determine the film thickness. An example SEM image of a $120$ nm thick film deposited via HiPIMS at $10\%$ nitrogen concentration is shown in figure 2.
@@
The resulting deposition rates are shown in figure 3. After these were determined, $120$ nm thick films were deposited at different nitrogen gas concentrations using both DC sputtering and HiPIMS, and the $T_{\mathrm{C}}$ of each film was measured. Here we define $T_{\mathrm{C}}$ as the temperature at which the resistance drops to half of its normal-state value at 18 K. Figure 3 shows these results. The nitrogen concentration yielding the highest $T_{\mathrm{C}}$ was 15% for DC sputtering and 20% for HiPIMS.
Figure 3: (a) $T_{\mathrm{C}}$ of $120$ nm thick NbN films deposited via HiPIMS and DC sputtering onto unheated oxidized silicon wafers with different nitrogen gas concentrations. (b) Deposition rate as a function of nitrogen gas concentration. DC sputtering was performed at $150$ W and HiPIMS was performed at a time-average power of $250$ W.
X-ray diffraction (XRD) was used to determine the lattice constant of films deposited via DC sputtering and HiPIMS at the optimal nitrogen concentration for each process. To increase the XRD signal, thicker films of $1$ $\mu$m were grown for this characterization. The results are shown in figure 4. Peaks consistent with a cubic crystal structure and corresponding to Miller indices ($hkl$) = (111) and (200) are observed. For a simple cubic structure, the lattice constant $a$ is related to the interplane spacing $d$ by $a=d\sqrt{h^{2}+k^{2}+l^{2}}$. The diffraction angle $\theta$ is related to $d$ via Bragg’s law, $n\lambda=2d\sin\theta$, where here we have first-order diffraction ($n=1$). The XRD system uses a Cu Kα x-ray source, which has a wavelength $\lambda=1.5406$.
For each XRD peak in figure 4, a Gaussian fit was performed to extract the peak center. This angle was then used to calculate the associated lattice constant $a$. The results are given in table 1. We find a lattice constant of $4.40$ for the HiPIMS film and a smaller lattice constant of $4.38$ for the DC sputtered film. Previous work has found that an increase in nitrogen concentration is associated with an increase in the lattice constant.Shy, Toth, and Somasundaram (1973); Kalal, Gupta, and Rawat (2021); Torche et al. (2003) The larger lattice constant in the HiIPMS film is consistent with the picture of the HiPIMS process enabling us to get closer to optimal stoichiometry before beginning to form the $\epsilon$-NbN crystal phase that reduces $T_{\mathrm{C}}$.
Figure 4: X-ray diffraction of $1$ $\mu$m thick films deposited via DC sputtering and HiPIMS using the optimal nitrogen concentration for each process, which was 15% for DC sputtering and 20% for HiIPMS.
@@
$2\theta$ peak lattice constant $2\theta$ peak lattice constant
$35.323$ $4.3976$ $35.493$ $4.3772$
$41.027$ $4.3964$ $41.141$ $4.3847$
Figure 5: $T_{\mathrm{C}}$ of NbN films deposited via HiPIMS and DC sputtering as a function of film thickness. Films were deposited onto unheated silicon substrates with a native oxide using the optimal nitrogen concentration for each deposition technique.
Using the optimal nitrogen concentration for the DC and HiPIMS processes respectively, a set of films of varying thickness between $6$ nm and $120$ nm were deposited to explore the dependence of $T_{\mathrm{C}}$ on the film thickness. The results are shown in figure 5. We see that the HiPIMS films have a higher $T_{\mathrm{C}}$ than the DC sputtered films at all thicknesses studied, with the difference becoming larger for the thinnest samples.
@@ IV Optimization of HiPIMS Films
To study the potential for improvement of the $T_{\mathrm{C}}$ of our HiPIMS films, we deposited a set of films of different thicknesses on top of a $20$ nm AlN buffer layer, and on top of a heated substrate with a $40$ nm AlN buffer layer. The AlN layer was deposited as described previously, and the the NbN was then deposited in the same deposition system without breaking vacuum. All NbN films were deposited using HiPIMS with a $20$% nitrogen gas concentration. The resulting critical temperatures are shown in figure 6. As has been seen in previous work with DC sputtering, the use of the AlN buffer layer increased $T_{\mathrm{C}}$. This is attributed to lattice matching between the AlN and cubic NbN.Shiino et al. (2010); Rhazi et al. (2021)
@@
First, we deposited NbN onto bare substrates (Si with a native oxide and no AlN buffer layer) that were heated to $350$ C. The resistance versus temperature curves of the resulting films showed a reduced and significantly broadened superconducting transition ($\Delta T_{\mathrm{C}}>1$ K). We believe that this resulted from an unwanted chemical interaction between the substrate and the NbN enabled by the combination of thermal energy and the high kinetic energy produced by the HiPIMS pulses. We found that this could be prevented and sharp transitions could be obtained by first depositing the AlN layer at room temperature, then heating the substrate to $350$ C, and then depositing the NbN. For these samples, the thickness of the AlN buffer layer was increased to $40$ nm to ensure that it was effective at inhibiting the interaction between the NbN and the underlying substrate. With the buffer layer, the heated NbN films showed a further increase in $T_{\mathrm{C}}$, as seen in figure 6. This is consistent with the expectation that a higher substrate temperature results in greater adatom mobility during the deposition process, which improves crystallization as local nitrogen excesses and deficits in the lattice can diffuse out and long-range crystalline order can develop.Kalal, Gupta, and Rawat (2021); Thakoor et al. (1985)
@@
In each case, the HiPIMS process produced films with a lower sheet resistance and a higher $T_{\mathrm{C}}$ than DC sputtering. Interestingly, and in contrast to previous results,Shiino et al. (2010) we observe an increase in the sheet resistance when the film is deposited on an AlN buffer layer as compared to the bare substrate. We speculate that this may arise from the surface roughness of the sputtered AlN. If this is the case, one would expect to see this effect reduced in thicker NbN films. To test this, we deposited two $50$ nm thick NbN films via HiPIMS at $20$% nitrogen gas concentration, one with a $40$ nm AlN buffer layer and one without the buffer layer. These films both had the same sheet resistance, $37$ $\Omega$/sq, at room temperature. This is consistent with the idea that the higher sheet resistance of the $10$ nm NbN films on AlN is related to AlN surface roughness. In future work, it may be useful to explore approaches for improving the surface roughness of the sputtered AlN buffer layer.
Figure 6: $T_{\mathrm{C}}$ as a function of thickness of NbN films deposited via HiPIMS with 20% nitrogen concentration. Deposition onto unheated silicon substrates with a native oxide is compared to films deposited on the same substrate with the addition of a $20$ nm thick AlN buffer layer, as well as films with a $40$ nm thick AlN buffer layer that were heated to $350$ C for the NbN deposition.
@@
Deposition type Sheet resistance at 296 K Sheet resistance at 18 K Critical temperature
DC $652$ $\Omega$/sq $908$ $\Omega$/sq $5.90$ K
HiPIMS $281$ $\Omega$/sq $328$ $\Omega$/sq $9.30$ K
DC with AlN buffer 921 $\Omega$/sq 1,613 $\Omega$/sq 7.25 K
HiPIMS with AlN buffer $346$ $\Omega$/sq $419$ $\Omega$/sq $11.41$ K
heated DC with AlN buffer 691 $\Omega$/sq 1,129 $\Omega$/sq 8.90 K
heated HiPIMS with AlN buffer $323$ $\Omega$/sq 391 $\Omega$/sq $12.55$ K

### v2

@@ Abstract
We report a systematic comparison of niobium nitride thin films deposited on oxidized silicon substrates by reactive DC magnetron sputtering and reactive high-power impulse magnetron sputtering (HiPIMS). After determining the nitrogen gas concentration that produces the highest superconducting critical temperature for each process, we characterize the dependence of the critical temperature on film thickness. The optimal nitrogen concentration is higher for HiPIMS than for DC sputtering, and HiPIMS produces higher critical temperatures for all thicknesses studied. We attribute this to the HiPIMS process enabling the films to get closer to optimal stoichiometry before beginning to form a hexagonal crystal phase that reduces the critical temperature, along with the extra kinetic energy in the HiPIMS process improving crystallinity. We also study the ability to increase the critical temperature of the HiPIMS films through the use of an aluminum nitride buffer layer and substrate heating.
@@ I Introduction
Niobium nitride (NbN) is a refractory metal with a superconducting critical temperature ($T_{\mathrm{C}}$) as high as $\approx 16$ K.Shy, Toth, and Somasundaram (1973) NbN thin films are utilized in a wide range of superconducting device applications, including photodetectors such as hot electron bolometersGousev et al. (1994) and superconducting nanowire single-photon detectors;Holzman and Ivry (2019) and Josephson junction devices such as SIS heterodyne mixers,McGrath et al. (1991) RSFQ digital logic,Shao et al. (2024) and superconducting qubits.Kim et al. (2021) Due to their high kinetic inductivity,Annunziata et al. (2010) NbN thin films are of interest in applications such as slow-wave transmission lines,Santavicca et al. (2021) tunable couplers,Colangelo et al. (2021) and travelling wave parametric amplifiers.Adamyan et al. (2016) For many of these applications, very thin films ($\lesssim 10$ nm) are required, and hence a number of groups have explored different techniques for producing high-quality ultra-thin NbN films.Ilin et al. (2008); Espiau de Lamaëstre, Odier, and Villégier (2007); Iovan et al. (2023); Chockalingam et al. (2008); Wright, Xing, and Jena (2023); Schneider et al. (2009); Shiino et al. (2010); Rhazi et al. (2021)
Several crystal phases of NbN can occur under normal conditions. The cubic NaCl-like $\delta$-phase has the highest $T_{\mathrm{C}}$ and occurs when a near-stoichiometric reaction between nitrogen and niobium takes place.Shy, Toth, and Somasundaram (1973) At a $1{:}1$ stoichiometry, however, the most stable crystal phase is the hexagonal $\epsilon$-phase, which has a lower $T_{\mathrm{C}}$. On the other hand, reducing the nitrogen concentration can lead to the formation of tetragonal Nb4N3 or hexagonal Nb2N, which also have a lower $T_{\mathrm{C}}$.Benkahoul et al. (2004); Wright, Xing, and Jena (2023) Optimizing the nitrogen concentration is thus important for maximizing the $T_{\mathrm{C}}$ of NbN thin films.
@@
Heating the substrate can improve the crystallinity of the film, and single-domain epitaxial films can be produced using an appropriate substrate at sufficiently high temperature.Kim et al. (2021); Wright, Xing, and Jena (2023); Iovan et al. (2023); Chockalingam et al. (2008) These high temperatures, however, are not compatible with some materials and fabrication processes, including lift-off processes and standard CMOS processes. Single-domain epitaxial films also have a lower normal-state resistivity and hence a lower kinetic inductivity than polycrystalline films,Annunziata et al. (2010) and as a result they may not be desirable for applications that utilize the high kinetic inductivity of NbN.
@@ II Experimental Parameters
NbN films were deposited using a Nb target in a reactive process with nitrogen gas. The Nb target was specified as 99.95% pure (excluding tantalum). Ultra high purity argon gas (99.9999% purity) and ultra high purity nitrogen gas (99.999% purity) were introduced into the chamber via separate mass flow controllers. The chamber was pumped to $\lesssim 5\times 10^{-7}$ torr prior to deposition. During all depositions, a gate valve is used to throttle the turbomolecular pump and the mass flow controllers set the gas flow rates to achieve a chamber pressure of $3.0\times 10^{-3}$ torr. The nitrogen gas concentration is specified as a percentage of the argon gas flow rate (not as a percentage of the total gas flow). For example, an argon flow rate of $12.5$ sccm and a nitrogen flow rate of $2.5$ sccm would be specified as a $20$% nitrogen gas concentration.
Depositions were performed using both DC sputtering and HiPIMS. The HiPIMS depositions utilized a unipolar DC power supply connected to an Impulse $2$ kW pulsed power module from Starfire Industries. This module allows for a user-configurable voltage pulse height, width, and repetition rate. All HiPIMS depositions described here used a $1$ kHz repetition rate and a $100$ $\mu$s pulse width. The HiPIMS was performed in a power-limited mode set to a time-average power of $250$ W; this determined the resulting voltage pulse height, which had a value of approximately $480$ V. The main pulse was followed by a positive kick pulse with an amplitude of approximately $+50$ V and a duration of $50$ $\mu$s. This kick pulse has been shown to increase the deposition rate by driving ionized target material to the sample, as the high peak power of the HiPIMS process can result in a higher ionization fraction of the target material.Gaines (2019); Kalal et al. (2022) The kick pulse has also been shown to increase the average ion kinetic energy.Keraudya et al. (2019) An example of the resulting voltage and current pulses is shown in figure 1. The peak power, which occurs at the maximum of the current pulse, is approximately $3$ kW.
@@
Electrical measurements of the NbN films were performed in a dipper-style liquid helium cryostat with the samples inside a vacuum can. The system has a $1$ K pot and a base temperature of approximately $1.5$ K. The films were mounted directly onto a copper sample holder using GE varnish, and the copper sample holder was attached to a copper block connected to the $1$ K pot. A thermometer and resistive heater are also mounted to the copper block. Ten Manganin wires (36 AWG) connect between the sample space and room temperature, with each pair of wires having a 2-wire lead resistance of $120$ $\Omega$. Aluminum wirebonds provide the electrical connection between the film and sample holder. The device resistance is measured with a lock-in amplifier using an AC excitation current with an RMS amplitude of $1$ $\mu$A, and resistance versus temperature curves are taken by first cooling the device below $T_{\mathrm{C}}$, then using the heater to heat the sample up to $18$ K, and then monitoring the device voltage as the heater voltage ramps down to gradually cool the sample through its superconducting transition. The rate of temperature decrease was approximately $0.3$ K per minute, ensuring negligible thermal lag between the sample and thermometer.
@@ III Comparison of DC and HiPIMS Films
NbN films were deposited onto unheated substrates at varying nitrogen gas concentrations using both DC sputtering and HiPIMS. The deposition rate varies with the nitrogen gas concentration, so the deposition rate had to be determined for each nitrogen concentration. This was done by performing a $900$ s deposition, scribing and snapping the resulting sample, and imaging the sample’s cross-section in a scanning electron microscope (SEM) to determine the film thickness. An example SEM image of a $120$ nm thick film deposited via HiPIMS at $10\%$ nitrogen gas concentration is shown in figure 2.
@@
The resulting deposition rates are shown in figure 3b. After these were determined, $120$ nm thick films were deposited at different nitrogen gas concentrations using both DC sputtering and HiPIMS, and the $T_{\mathrm{C}}$ of each film was measured. Here we define $T_{\mathrm{C}}$ as the temperature at which the resistance drops to half of its value at $18$ K. Figure 3a shows these results. The nitrogen concentration yielding the highest $T_{\mathrm{C}}$ was 15% for DC sputtering and 20% for HiPIMS. The higher nitrogen gas concentration for the optimal HiPIMS film suggests that the HiPIMS process enables us to get closer to $1{:}1$ stoichiometry before beginning to form the $\epsilon$-NbN crystal phase that reduces $T_{\mathrm{C}}$.
Figure 3: (a) $T_{\mathrm{C}}$ of $120$ nm thick NbN films deposited via HiPIMS and DC sputtering onto unheated oxidized Si wafers with different nitrogen gas concentrations. (b) Deposition rate as a function of nitrogen gas concentration. DC sputtering was performed at $150$ W and HiPIMS was performed at a time-average power of $250$ W.
To gain further insight into the film composition and microsctructure, x-ray diffraction (XRD) $\theta$-$2\theta$ scans were performed with a Shimadzu XRD-6100 on films deposited via DC sputtering and HiPIMS at the optimal nitrogen concentration for each process. To increase the XRD signal, thicker films of $1$ µm were grown for this characterization. $2\theta$ peaks were observed at approximately $35^{\circ}$ and $41^{\circ}$, as shown in figure 4. Each measured spectrum has been normalized to the height of the larger peak. The observed peaks are consistent with Miller indices ($hkl$) = (111) and (200) for a cubic crystal structure. For a simple cubic structure, the lattice constant $a$ is related to the interplane spacing $d$ by $a=d\sqrt{h^{2}+k^{2}+l^{2}}$. The diffraction angle $\theta$ is related to $d$ via Bragg’s law, $n\lambda=2d\sin\theta$, with $n=1$. The XRD system uses a Cu Kα x-ray source with a wavelength $\lambda=1.5418$.Harrington and Santiso (2021) Each measured peak was fit to a pseudo-Voigt function, the weighted sum of a Gaussian and a Lorentzian. The peak center from the fit was then used to calculate the lattice constant $a$ using the specified Miller indices. The results are summarized in table 1. Also included in the table are the values of the full-width at half-maximum (FWHM) of each peak.
The HiPIMS film yielded a lattice constant of $4.400$-$4.401$ and the DC sputtered film yielded a lattice constant of $4.382$-$4.388$. Previous work has shown that an increase in nitrogen concentration is associated with an increase in the lattice constant.Shy, Toth, and Somasundaram (1973); Kalal, Gupta, and Rawat (2021); Torche et al. (2003) The larger lattice constant in the HiPIMS film thus confirms that this film has higher nitrogen content than the DC sputtered film, consistent with the picture of the HiPIMS process enabling us to get closer to optimal stoichiometry before beginning to form the $\epsilon$-NbN crystal phase that reduces $T_{\mathrm{C}}$.
@@
$2\theta$ peak FHWM lattice constant $2\theta$ peak FWHM lattice constant
$35.324^{\circ}$ $0.237^{\circ}$ $4.401$ $35.483^{\circ}$ $0.341^{\circ}$ $4.382$
$41.027^{\circ}$ $0.441^{\circ}$ $4.400$ $41.141^{\circ}$ $0.716^{\circ}$ $4.388$
Figure 4: X-ray diffraction spectrum showing measured peaks (markers) along with pseudo-Voigt fits (dashed lines). Samples are $1$ $\mu$m thick films deposited via HiPIMS and DC sputtering using the optimal nitrogen concentration for each process. (a) shows the (111) peaks and (b) shows the (200) peaks, and and each spectrum has been normalized to the height of the (111) peak.
The peak width contains contributions from crystallite size, film disorder (including strain), and instrument resolution. In the pseudo-Voigt fit, the width of the Lorenzian component is attributed to crystallite size, while the width of the Gaussian component is attributed to nonuniform strain.Harrington and Santiso (2021) For all our peak fits, the Lorentzian component dominated – the pseudo-Voigt fit was nearly identical to a Lorentzian fit – suggesting that the dominant source of intrinsic peak broadening is crystallite size. Assuming that the sample broadening is due only to finite crystallite size, we can determine the volume-average crystallite size $L_{V}$ using the Scherrer formula, $L_{V}=K\lambda/(\beta_{s}\cos\theta)$, where $K$ is the unitless structure factor, which for simplicity we assume to be $0.94$ for spherical crystallites, $\beta_{s}$ is the intrinsic sample broadening (in radians), and $\theta$ is the peak angle. The sample broadening is found from $\beta_{s}=\sqrt{\beta_{m}^{2}-\beta_{i}^{2}}$, where $\beta_{m}$ is the measured FWHM of each peak and $\beta_{i}$ is the contribution from the instrument resolution. The instrument resolution was quantified by measuring a Si powder reference, whose (111) peak was centered at $2\theta=28.406^{\circ}$ with a FWHM of $\beta_{i}=0.129^{\circ}$. Using this approach, for the (111) peak of the HiPIMS film we get $L_{V}=48$ nm, and for the (111) peak of the DC sputtered film we get $L_{V}=29$ nm. We note that the (200) peaks are broader and yield smaller crystallite sizes than the (111) peaks, although the trend of the HiPIMS film having a larger crystallite size still holds. This discrepancy could be related to asymmetrical crystallites, as each peak width reflects the crystallite size in that particular (hkl) direction, and it could also be related to other sources of peak broadening not captured by the Scherrer equation. Nonetheless, the narrower peak widths in the XRD spectrum of the HiPIMS film indicate that the HiPIMS process produces films with a higher degree of crystallinity than DC sputtering.
Figure 5: $T_{\mathrm{C}}$ of NbN films deposited via HiPIMS and DC sputtering as a function of film thickness. Error bars show the width of the transition, defined as 90% to 10% of the normal-state resistance at $18$ K. Films were deposited onto unheated Si substrates with a native oxide using the optimal nitrogen concentration for each deposition technique.
Figure 6: Resistance versus temperature curve from a $10$ nm thick NbN film deposited via HiPIMS with $20\%$ nitrogen gas concentration onto an unheated substrate. The resistance has been normalized to the value at $18$ K, and $120$ $\Omega$ of lead resistance was subtracted before normalization. $T_{\mathrm{C}}$ is defined as the temperature at which the resistance is half the value at $18$ K. The transition width $\Delta T_{\mathrm{C}}=T_{high}-T_{low}$, where $T_{high}$ is the temperature at which the resistance is $90\%$ of the value at $18$ K and $T_{low}$ is the temperature at which the resistance is $10\%$ of the value at $18$ K.
Using the optimal nitrogen concentrations for the DC and HiPIMS processes, a set of films of varying thickness between $6$ nm and $120$ nm were deposited in order to explore the dependence of $T_{\mathrm{C}}$ on the film thickness. The results are shown in figure 5. As before, $T_{\mathrm{C}}$ is defined as the temperature at which the resistance drops to half the value at $18$ K. We see that the HiPIMS films have a higher $T_{\mathrm{C}}$ than the DC sputtered films at all thicknesses studied, with the difference becoming larger for the thinnest samples. The error bars in figure 5 show the transition width $\Delta T_{\mathrm{C}}=T_{high}-T_{low}$, where $T_{high}$ is the temperature at which the resistance is $90\%$ of the value at $18$ K and $T_{low}$ is the temperature at which the resistance is $10\%$ of the value at $18$ K (after subtraction of the lead resistance). The transition width broadens as the film thickness is decreased and develops a noticeably asymmetry, with $(T_{high}-T_{\mathrm{C}})>(T_{\mathrm{C}}-T_{low})$, as seen in the example resistance versus temperature curve shown in figure 6. This asymmetry is attributed to fluctuations of the superconducting order parameter above $T_{\mathrm{C}}$, which are enhanced in systems with low dimensionality and strong disorder.Aslamasov and Larkin (1968); Baeva et al. (2024)
@@ IV Optimization of HiPIMS Films
To study the potential for improvement of the $T_{\mathrm{C}}$ of our HiPIMS films, we deposited a set of films of different thicknesses on top of a $20$ nm AlN buffer layer, and on top of a heated substrate with a $40$ nm AlN buffer layer. The AlN layer was deposited as described previously, and then the NbN was deposited in the same deposition system without breaking vacuum. All NbN films were deposited using HiPIMS with a $20$% nitrogen gas concentration. The resulting critical temperatures are shown in figure 7, with the error bars showing $\Delta T_{\mathrm{C}}$. As has been seen in previous work with DC sputtering, the use of the AlN buffer layer increased $T_{\mathrm{C}}$.Shiino et al. (2010); Rhazi et al. (2021) This is attributed to lattice matching between the c-plane of hexagonal AlN and the (111) plane of cubic NbN.
@@
First, we deposited NbN onto bare substrates (Si with a native oxide and no AlN buffer layer) that were heated to $350$ C. The resistance versus temperature curves of the resulting films showed a reduced and significantly broadened superconducting transition ($\Delta T_{\mathrm{C}}>1$ K for relatively thick films). Films deposited with DC sputtering onto substrates heated to the same temperature did not show this degradation of $T_{\mathrm{C}}$. Thus we believe that the combination of thermal energy and the high kinetic energy produced by the HiPIMS pulses enables an unwanted chemical interaction between the substrate and the NbN. We found that this could be prevented, and relatively sharp transitions could be obtained, by first depositing the AlN layer at room temperature, then heating the substrate to $350$ C, and then depositing the NbN. For these samples, the thickness of the AlN buffer layer was increased to $40$ nm to ensure that it was effective at inhibiting the interaction between the NbN and the underlying substrate. With the buffer layer, the heated NbN films showed a further increase in $T_{\mathrm{C}}$, as seen in figure 7. This is consistent with the expectation that a higher substrate temperature results in greater adatom mobility during the deposition process, which improves crystallization as local nitrogen excesses and deficits in the lattice can diffuse out and long-range crystalline order can develop.Kalal, Gupta, and Rawat (2021)
Figure 7: $T_{\mathrm{C}}$ as a function of thickness of NbN films deposited via HiPIMS with 20% nitrogen concentration. Deposition onto unheated Si substrates with a native oxide is compared to films deposited on the same substrate with the addition of a $20$ nm thick AlN buffer layer, as well as films with a $40$ nm thick AlN buffer layer that were heated to $350$ C for the NbN deposition. Error bars show the width of the transition, defined as 90% to 10% of the normal-state resistance at $18$ K.
@@
Deposition type Sheet resistance at 296 K Sheet resistance at 18 K $T_{\mathrm{C}}$ $\Delta T_{\mathrm{C}}$
DC $652$ $\Omega$/sq $908$ $\Omega$/sq $5.90$ K $1.70$ K
HiPIMS $281$ $\Omega$/sq $328$ $\Omega$/sq $9.36$ K $0.89$ K
DC with AlN buffer 921 $\Omega$/sq 1,613 $\Omega$/sq 7.29 K $2.72$ K
HiPIMS with AlN buffer $346$ $\Omega$/sq $419$ $\Omega$/sq $11.42$ K $0.97$ K
heated DC with AlN buffer 691 $\Omega$/sq 1,129 $\Omega$/sq 8.96 K $2.32$ K
heated HiPIMS with AlN buffer $323$ $\Omega$/sq 391 $\Omega$/sq $12.44$ K $0.98$ K
In each case, the HiPIMS process produced films with a lower sheet resistance and higher $T_{\mathrm{C}}$ than DC sputtering. Interestingly, and in contrast to previous results,Shiino et al. (2010) we observe an increase in the sheet resistance when the film is deposited on an AlN buffer layer as compared to the bare substrate. To determine if this is related to the surface roughness of the AlN, a $20$ nm thick AlN film was characterized via atomic force microscopy. The resulting root mean square roughness of $0.4$ nm is not likely to significantly affect the resistivity of $10$ nm thick NbN films. It may be that the increased resistivity is related to film strain, although further studies would be required to test this hypothesis.